OpenStudy (loser66):

Find all solutions of \(980x \equiv 1500 (mod 1600)\) Please , help

OpenStudy (loser66):

When solving Diophantine, I got the final equation is \(980 (-31) - 1600(-19) =20\) Multiple both sides by 75, I got \(980 (-2325) - 1600 (-1425) = 1500\) But I don' t know how to get the answer from back of the book x = 75 + 80 k

OpenStudy (loser66):

@ganeshie8

ganeshie8 (ganeshie8):

You could reduce the amount of work by dividing out any common factors

ganeshie8 (ganeshie8):

\(980x \equiv 1500 \pmod {1600}\) divide 20 through out and get \(49x \equiv 75 \pmod {80}\)

ganeshie8 (ganeshie8):

both congruences are "equivalent"

OpenStudy (loser66):

Can I? I didn't see any theorem or corollary allows me to do that. However, suppose I can, then?

ganeshie8 (ganeshie8):

You don't really need a huge theorem to do that, just write out what a congruence means and use divisibility argument

ganeshie8 (ganeshie8):

\(980x \equiv 1500 \pmod {1600} \iff 1600 \mid (980x-1500)\) yes ?

OpenStudy (loser66):

yes

ganeshie8 (ganeshie8):

\(980x \equiv 1500 \pmod {1600} \iff 1600 \mid (980x-1500)\) \(\iff 80*20 \mid 20(49x-75)\) yes ?

OpenStudy (loser66):

yes

ganeshie8 (ganeshie8):

can we say \(ab \mid ac \iff b \mid c\) ?

OpenStudy (loser66):

yes

ganeshie8 (ganeshie8):

\(980x \equiv 1500 \pmod {1600} \iff 1600 \mid (980x-1500)\) \(\iff 80*20 \mid 20(49x-75)\) \(\iff 80 \mid (49x-75)\) yes ?

OpenStudy (loser66):

yes, :)

ganeshie8 (ganeshie8):

good :)

OpenStudy (loser66):

Now, I solve this equivalent and the result is the same with original one, right?

ganeshie8 (ganeshie8):

Yep. your original equation is like 2y + 2x = 2 your new equation is like y + x = 1 they have same solutions because they are identical equations

OpenStudy (loser66):

Thank you so much, let me try.

ganeshie8 (ganeshie8):

you may double check with wolfram http://www.wolframalpha.com/input/?i=solve+49x%3D75+mod+80

OpenStudy (loser66):

Question : if we deduce the congruence like above, we miss one solution. Problem: Solve \(9x\equiv 6~(mod~15)\) 1) Solve as it is: 9x -15y = 6 gcd (9,15) =3 15 = 9 + 6 9 = 6 + 3 6 = (2) 3 + 0 Hence 3 = 9 - 6 = 9 - ( 15 -9) = (2) 9 - 15 Hence 6 = (4) 9 -(2) 15 \(x_0 = 4 \\x = 4 + \dfrac{15}{3}t; t={0,1,2}\) Hence \(x = \overline 4; x= \overline 9; x =\overline {14}\) 2) deduce \(9x\equiv 6 ~(mod~15)\\3x\equiv 2~(mod~5)\) That is, we need find 3x -5y = 2 gcd (3,5) =1 5 = 3 + 2 --> 2 = 5 - 3 3 = 2 + 1--> 1 = 3 -2 1 = 3- (5 - 3) = (2) 3 -5 hence 3(4) - 5(2) = 2 \(x_0 = 4: x = 4 +5t: t =\{0,1\}\) Then \(x = \overline 4 ~~or~~ x=\overline 9\). We miss \(x =\overline {14}\) @ganeshie8

ganeshie8 (ganeshie8):

what makes you think we're missign x=14 ?

ganeshie8 (ganeshie8):

@Loser66

OpenStudy (loser66):

if it is mod 5, then x =14 = 4 , but we have to solve for mod 15, hence x = 14 must be one of the solutions. If we make the equivalent congruent as above and solve it, we don't have x =14 as it is in original one.

ganeshie8 (ganeshie8):

I see what you're thinking. For the congruence, \(9x\equiv 6 ~\pmod{15}\), the solution is \(x\equiv 4\pmod{5}\). How do you express that solution in parameter form ?

ganeshie8 (ganeshie8):

@Loser66

OpenStudy (loser66):

@ganeshie8

OpenStudy (loser66):

I don't get your question, x =4 mod 5 , \(x = 4 + 5t\)

ganeshie8 (ganeshie8):

Yes, what integers does that solution include ? Maybe write out first few

OpenStudy (loser66):

1

OpenStudy (loser66):

0

ganeshie8 (ganeshie8):

don't wry about the mod, just write out the integers, x, that satisfy \(9x\equiv 6\pmod{15}\)

OpenStudy (loser66):

x = 4, 9, 14

ganeshie8 (ganeshie8):

why only 3 ? doesn't -1 satisfy the congruence ?

OpenStudy (loser66):

because -1 =4 mod 5

ganeshie8 (ganeshie8):

I said, forget about mod

OpenStudy (loser66):

oh, because we have the congruent class, not a concrete number.

OpenStudy (loser66):

and -1 belongs to {14 bar}

ganeshie8 (ganeshie8):

write out first few integers, x, that satisfy the given congruence

ganeshie8 (ganeshie8):

I'm just asking you to write out first few integer, x, that satisfy the given congruence. That is a simple question. You don't really need to use mod.

OpenStudy (loser66):

-1,4,6

OpenStudy (loser66):

is it not that when we write out something, we need a mod to calculate ? like if you say 9x = 6 , I have to know on which mod to write out some solutions. If I don't have any mode, how can I know them since 9x =6 mod 2 differ from mod 3, mod 4, mod 5 and so on. ... I hate to be dummy like this but I do not get even a "simple question" like what you think.

ganeshie8 (ganeshie8):

we don't "need" another congruence to write out solutions to a congruence.

ganeshie8 (ganeshie8):

Let me write out few solutions to the congruence \(9x\equiv 6\pmod{15}\) : the integers \(x = 4,9,14,-1,-6\) satisfy above congruence

OpenStudy (loser66):

I think you should teach me all. I can learn quickly. It is better than you ask something you assume that I knew but I didn't. :)

ganeshie8 (ganeshie8):

I am pretty sure you knew the theory. It's just that you had never used congruences much, you're finding it difficult to interpret them properly...

ganeshie8 (ganeshie8):

As you can see, every element in the set, \(\{4+5t : t\in \mathbb{Z}\}\), is a solution to the congruence \(9x\equiv 6\pmod{15}\)

OpenStudy (loser66):

Sanity is back

ganeshie8 (ganeshie8):

But that set form looks ugly when to do arithmetic, so we express the solutions in congruence form instead : \(x\equiv 4\pmod{5}\)

ganeshie8 (ganeshie8):

When we say, \(x\equiv 4\pmod{5}\) is the solution, we're saying that the infinite set \(\{4+5t : t\in \mathbb{Z}\}\) is the solution.

OpenStudy (loser66):

but in class, it is not all t in Z, they are remainder of division by gcd ( a, b) on ax + by = c

ganeshie8 (ganeshie8):

I said, forget about mod

OpenStudy (loser66):

ok

ganeshie8 (ganeshie8):

the question is about finding the solutions to given congruence not about how we express them

ganeshie8 (ganeshie8):

But if somebody asks you to list out the "incongruent" solutions in mod 15, what would you list ?

ganeshie8 (ganeshie8):

then you can say a set of "incongruent" solutions in mod 14 is : \(\{4,9,14\}\)

ganeshie8 (ganeshie8):

Somebody else comes up and asks you to list out the "incongruent" solutions in mod 5, then you would give them a set with one element : \(\{4\}\)

ganeshie8 (ganeshie8):

Somebody else comes up and asks you to list out the "incongruent" solutions in mod 10, then you would give them a set with one element : \(\{4,9\}\)

ganeshie8 (ganeshie8):

As you can see all those answers are same. They all give the set \(\{4+5t : t\in \mathbb{Z}\}\).

ganeshie8 (ganeshie8):

\(\{4,9,14\}\) in mod 15 is same as \(\{4\}\) in mod 5 is same as \(\{4,9\}\) in mod 10 is same as \(\{4+5t : t\in \mathbb{Z}\}\)

ganeshie8 (ganeshie8):

they all generate the same integers as solutions. they are just different but equivalent ways to express the solution to congruence \(9x\equiv 6\pmod{15}\)

ganeshie8 (ganeshie8):

It doesn't matter how you express the final solution if the textbook doesn't ask you explicitly

ganeshie8 (ganeshie8):

Ofcourse simply saying \(x\equiv 4\pmod{5}\) is the solution looks better. "Congruence" is just a short hand "notation" which makes expressing divisibility statements compact. Nothing more.

OpenStudy (loser66):

It is new to me. I have to digest it completely. Thanks a lot.

ganeshie8 (ganeshie8):

np, just ask if something is not clear. i think we are messing with two formats of expressing the solutions : 1) all the solutions of a congruence 2) incongruent solutions under specific mod you don't "need" to give the solutions in second format. just know that we can simply write out the solution in set form